OpenStudy (anonymous):

A 12 C sample of chlorine gas at 1.5 atm is warmed up to 35 C. What is its new pressure in mmHg?

OpenStudy (anonymous):

PV=nRT Assuming volume and number of moles are constant which is probably a fine assumping because we can assume the gas is in a rigid container which does not receive of lose any gas molecules... (and R is a constant) P1 /T1 = P2/T2 P1 is 1.5 am, T1 = 12 C + 273 = 285 Kelvin (must do this in kelvin because of units) P2 is unknown, T2 = 35 + 273 = 308 Kelvin 1.5 / 285 = x / 308 x = 1.5 * 308/285 = 1.62 atm 1.62atm * (760mmHg/1atm) = 1232 mmHg